Baltimore Company's complete assets and liabilities are Accounts Receivable $800, Equipment $10,000, Accounts Payable $4,100, Prepaid Rent $2,000, Supplies $400, Bank Loan $1,550, and Tools $300. Baltimore's total liabilities are: (All account balances are normal.)

Baltimore Company's complete assets and liabilities are Accounts Receivable $2,550, Equipment $8,800, Accounts Payable $7,300, Prepaid Rent $1,650, Supplies $775, Bank Loan $4,400, and Tools $535. Baltimore's total equity is: (All account balances are normal.)
Answer #1

a) Total liabilities

Account payable 4100
Bank loan 1550
Total liabilities 5650

b) Assets = 2550+8800+1650+775+535 = 14310

Liabilities = 7300+4400 = 11700

Total equity = 14310-11700 = 2610